# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v1

@@ 3.1.2 Learning based imputation
Lin et al. propose a novel objective by integrating representation learning and data recovery into a unified information-theoretic framework termed COMPLETER [80] for clustering task. Specifically, COMPLETER propose to learn intra-modal consistency by maximizing mutual information and learn modal imputation by minimizing conditional entropy. Dong et al. proposed a deep Gaussian processes model to impute the missing modalities and further enhance the performance of downstream tasks [81]. Motivated by contrastive learning, Yang et al. introduces a multimodal imputation paradigm termed SURE for robust multimodal imputation [82]. The method treats complete multimodal pairs as positive and randomly sampling unpaired samples as negative. A noise-robust contrastive loss is employed to alleviate the influence of false negatives. Li et al. proposed a prototype-based multimodal imputation method, which incorporates a two-stream model with dual attention layers and contrastive learning mechanism to learn modal-specific prototypes and model the cross-modal relationship [83]. Considering incomplete multimodal inputs, the model performs modality imputation using prototypes learned from the incomplete multimodal training data and the sample-prototype relationship inherited from the observed modalities. Xu et al. [84] proposed a novel federated multi-view clustering method using an unsupervised technique to evaluate and refine imputation quality, efficiently handling various scenarios of incomplete multi-view data.
@@ 3.2.1 Latent representation and projection learning
Zhou et al. [87] integrated the latent representation learning and regression into a unified framework for Alzheimer’s disease diagnosis based on incomplete multimodal neuroimaging and genetic data. The motivation of the latent representation learning module for incomplete multimodal data is similar to model (4). However, a widely recognized limitation of the above methods based on model (4) is that they are only applicable to the case with bi-modal data. To address this limitation, weighted matrix factorization is introduced, where doubly aligned incomplete multimodal clustering [88], one-pass incomplete multimodal clustering [89], and localized sparse incomplete multimodal clustering [90] are representative methods. These models generally impose the location information of missing instances as the weighted matrices with binary values on the matrix factorization model to eliminate the loss yet reduce the negative influence of the missing instances. The basic model of the weighted matrix factorization based methods can be formulated as follows:
@@
Compared with the basic model (4), the methods based on model (5) are more flexible and can be easily extended to the data with more than two modalities and arbitrary incomplete cases. Besides of these weighted matrix factorization models, a generalized incomplete multimodal clustering model proposed to learn a complete common representation from the incomplete modal-specific representations with a partially aligned learning term [91]. Specifically, such a partially aligned learning term can be regarded as a variant of the weighted common representation learning. To prevent the model from overly focusing on certain modalities with larger dimensions, projective incomplete multimodal clustering transforms the matrix factorization into a variant of the weighted projection learning model to obtain the consensus representation from the incomplete multimodal data [92]. For the incomplete multimodal and incomplete multi-label classification task, a variant of the weighted projection learning model is developed to explore the information of available modalities [93]. In terms of infant neuroscience, Zhang et al. propose to explore incomplete multimodal data by constructing a latent representation to capture the complementary information from incomplete time-points [94]. In summary, the above methods focus on exploring the available information of the observed modalities by introducing some techniques like the weighting mechanism to eliminate the loss of those missing modalities.
@@ 3.2.2 Graph learning based
Graph learning based method is very popular in conventional multimodal clustering. This technique is also widely considered and integrated into other methods like matrix factorization based and deep learning based methods, to explore the structure information of data so as to enhance the performance. For incomplete multimodal learning, the main challenges to apply the graph learning based approach are two folds: (1) How to obtain the intrinsic structure relationships among the available instances in each modality? (2) How to fully explore these incomplete relationships? For example, for the incomplete multimodal clustering tasks, one of the major challenges is to obtain the consensus complete graph or representation from these incomplete modal-specific relationships associated with all incomplete modalities. Specifically, we can formulate this issue as the following simple framework:
@@
where ${{{\bar{X}}^{\left(m\right)}}}$ denotes the set of the available instances in the $m$-th modality. ${{S^{\left(m\right)}}}$ represents the modal-specific graph learned from the $m$-th modal data ${{{\bar{X}}^{\left(m\right)}}}$. ${{W^{\left(m\right)}}}$ is the modal-available or missing indicator matrix as defined in model (5). Taking the incomplete multimodal clustering task for example, the target of general graph learning based method is to learn the consensus graph or representation $Z$ from the incomplete data $\left\{{{{\bar{X}}^{\left(m\right)}}}\right\}_{m=1}^{l}$ or graph $\left\{{{S^{\left(m\right)}}}\right\}_{m=1}^{l}$, such that a consensus clustering result can be obtained by performing k-means or spectral clustering on $Z$.
As one representative work, incomplete multimodal spectral clustering [95] proposed a unified framework that integrates the adaptive collaborative representation based modal-specific graph learning and co-regularized consensus representation learning. The first term enables the method to adaptively capture the intrinsic relationships among available instances of each modality and the second term seeks to obtain the consensus complete representation by fully exploring such similarity relationship. For the data with only two modalities, Wu et al. [96] proposed a method based on structured graph learning, which seeks to learn two modal-specific graphs corresponding to the two modalities with a partially aligned constraint on the samples with complete modalities. Then, the method obtains the modal-specific clustering from the two graphs and adopts the best one-to-one map to obtain the final consensus clustering result. The work in [97] pre-constructs several modal-specific graphs from the available instances of every modality and then learned a consensus graph from these graphs via a weighted consensus graph learning model with the constraint of pre-constructed confident neighbor information. In summary, except for the imputation based graph methods, most graph based incomplete multimodal learning methods seek to exploit certain similarity information among available instances to discover the clustering structure and clustering result. In [98], for Alzheimer’s disease prediction, a unified linear regression based model is developed, in which the incomplete multimodal data are segmented into several complete subsets for model optimization. Similar to the above method, Liu et al. [99] segmented the incomplete multimodal data into several complete subsets and used the sparse representation algorithm to construct the sub-graphs from these subsets. Then a modal-aligned hypergraph learning model is developed to perform the Alzheimer’s disease diagnosis.
@@ 3.2.3 Kernel learning based
Kernel learning based methods transform the learning phase from the raw-feature space into the kernel space. Similar to the graph learning based methods, each element of the kernel matrix denotes a kind of the similarity or connection values of the corresponding two samples. For incomplete multimodal data, the constructed kernels w.r.t. different modalities will also be incomplete with missing rows and columns. Thus, the target of kernel learning based methods for incomplete multimodal data is to explore the information of these incomplete kernels. In this branch, most methods consider kernel imputation as we mentioned above like model (3) while relatively few works focus on exploiting certain available kernel information. Late fusion incomplete multimodal clustering [100] is one of the representative kernel imputation-free methods for incomplete multimodal data. The method does not need to impute the missing kernel elements but tries to simultaneously impute the missing representations corresponding to the missing modalities and learn the consensus representation in a joint framework. Furthermore, to reduce the low-quality imputation of the latent representations, Liu et al. [101] proposed an efficient and effective regularized incomplete multimodal clustering method (EERIMVC), which integrates the initialized consensus representation as prior knowledge. In this model, such prior representation is obtained by the conventional complete kernel algorithm with zero or average filling. Therefore, the quality of the initialized consensus representation will be an important factor to the performance. To break the gap between representation learning and label prediction, a one-stage incomplete multimodal clustering method is proposed [102]. Compared with EERIMVC, this method replaces the consensus representation with a clustering label matrix and an orthogonal matrix.
@@ 3.2.4 Deep learning based
Recent years, deep learning received lots of attention in incomplete multimodal learning. For instance, Zhao et al. [103] proposed a deep partial multimodal clustering network, which first uses deep encoders to extract the latent features of multimodal data and then adopts the graph regularized partial multimodal clustering model to learn the consensus representation from the incomplete bi-modal data. For data with more than two modalities, Wen et al. [104] proposed several deep incomplete multimodal clustering networks based on deep autoencoder framework, cognitive learning [105], and graph neural network [106], respectively. The objective function of these methods can be unified as the following formula:
@@
The above methods focus on the observed instances to train the network and can flexibly handle all kinds of incomplete data with more than two modalities by introducing the location matrix ${W^{\left(m\right)}}$ as prior weight in function $f$. In addition, a common technique adopted in these three methods is the weighted latent representation fusion, in which the location information of missing modalities is adopted as fusion weight to reduce the negative influence of missing modalities. Lately, Xu et al. [108] proposed an imputation-free and fusion-free deep network for incomplete multimodal data clustering. The method mines the multimodal cluster complementarity in the subspace and transforms such complementary information as supervised information to guide the network training. However, this method needs a part of samples with fully observed modalities for model training.
@@ 3.3 Discussion
To sum up, the ways to handle incomplete data can be commonly divided into two categories, imputation-based incomplete multimodal learning and imputation-free incomplete multimodal learning. Methods that focus on imputing incomplete modalities can be classified into two types: one indirectly recovers the missing modalities through graph or kernel completion, and the other directly fills in the original data. For imputation-free incomplete multimodal learning methods, their design philosophy focuses more on utilizing available modalities to explore information and understand multimodal data. However, both types of methods for handling incomplete multimodal data face deeper challenges. For example, the quality assessment of imputed instances for missing modalities is commonly overlooked. Besides, using prior missing information to impute unknown modalities is itself difficult to bridge the information gap and information imbalance issues caused by modal missing.
@@ 4 Balanced Multimodal Learning
Different modalities are tightly correlated, since they depict the same concept from different views. This property inspires the flourish of multimodal learning, where multiple modalities are integrated and expected to enhance the understanding of related events or objects. However, despite the natural cross-modal correlation, each modality is with distinct data sources and forms. For instance, audio data is typically represented as one-dimensional waveforms, while visual data consists of images composed of pixels. On the one hand, this discrepancy endows each modality with different properties, such as convergence speed, and then makes it is hard to process and learn all modalities well simultaneously, bringing difficulty to joint multimodal learning. On the other hand, this discrepancy also reflects in the quality of unimodal data. Even though all modalities depict the same concepts, their information related to the target event or objects are with different amount. For example, considering an audio-visual sample of label meeting, the visual data apparently show the visual content of the meeting, which is easy to be recognized (see Figure.1c). While the corresponding audio data is a noisy street car sound, which is hard to build correlation with the label meeting. Information amount of visual modality is clearly more than that of audio modality. Due to the greedy nature of deep neural networks [9], multimodal model would tend to rely on just high-quality modality with sufficient target-related information, while under-fitting the other modalities. To address these challenges and improve the efficacy of multimodal models, recent studies have focused on strategies to balance the differences between modalities and enhance the overall performance of the model.
Figure 3: The learning curves (error-rate) of audio model (A), video model (V), and the naive joint audio-video (AV) model on Kinetics dataset. Solid lines plot validation error while dashed lines show train error. Figure is from [7]
@@ 4.1 Property-discrepancy method
In most multimodal models, different modalities are commonly trained jointly with uniform learning objectives. While this joint training strategy has been generally proven to be effective in practice, there are still potential risk brought by the discrepancy in data of multiple modalities. The discrepancy of multimodal data in sources and forms result in that each modality has its own learning property. For example, as shown in Figure 3, in the action recognition dataset Kinetics [111], which contains audio and visual modalities, the audio data tends to require fewer training steps than the vision data before the model begins to be overfitting [112, 113]. The widely-used joint training multimodal framework often neglects this inherent learning property discrepancy of unimodal data, which can potentially lead to negative impacts on the model’s performance. More importantly, based on recent study [7], the jointly trained multimodal model even could be outperformed by the best unimodal counterpart. Thus, it is essential to consider discrepancy in learning properties across each modality to improve the performance of the model. To address this issue and balance discrepancy in learning property, several studies have been proposed from different perspective, including learning-objective-based method, optimization-based method and architecture-based method.
@@ 4.1.1 Learning-objective-based method
Since jointly trained multimodal frameworks often only have uniform loss, which integrate the information of all modalities, the learning of each modality is hard to be observed or controlled. The learning-objective-based methods add unimodal loss and make it possible to targetedly adjust unimodal training:
@@
where $\alpha_{mm}$ and $\alpha_{i}$ is the weight of loss function. $n$ is the number of modalities. Since the loss weight determines the influence of corresponding loss function in parameter update, its selection is essential for controlling unimodal training. Wang et al. [7] suggested to determine the optimal loss weights by measuring the proposed overfitting-to-generalization ratio, by which it could comprehensively consider and balance the discrepancy in convergence speed of each modality.
@@ 4.1.3 Architecture-based method
Besides the above architecture-agnostic methods which focus on learning objective and optimization, methods considering the discrepancy in learning property problem from the perspective of model architecture are also proposed. At the fusion stage, Xiao et al. [115] proposed to randomly drop the network pathway of the modality with faster convergence speed altogether with certain probability. In this way, the training of corresponding modality is slowed down and its learning dynamics is ensured to be more compatible with other unimodal counterpart. Bedsides the late fusion architecture, studies that utilize low-level cross-modal fusion to enhance multimodal joint learning are also proposed. Zhou et al[116] utilized central connections to link low-level features in both video and audio streams, capturing spatial (for video frames) and temporal information, thereby achieving high-level semantic representations and improving joint multimodal learning.
@@
Reference Category Application
[7] Property discrepancy Optimization-based Action classification
[114] Optimization-based Sentiment analysis
[113] Architecture-based Action classification
[116] Architecture-based Deepfake detection
[117] Quality discrepancy Learning-objective-based Violence Detection
[118] Learning-objective-based Action classification
[119] Learning-objective-based Product understanding
[120] Learning-objective-based Audio-visual fine-grained classification
[121] Learning-objective-based Action classification
[8, 122] Optimization-based Action classification, sentiment analysis
[123] Optimization-based Rumor detection
[124] Optimization-based Audio-visual video parsing
[125] Architecture-based Sentiment analysis
[126] Architecture-based Multimodal Conditional Image Synthesis
[127] Architecture-based RGB-T Semantic Segmentation
[9] Data-argumentation-based Muli-view classification, gesture classification
[128] Data-argumentation-based Emotion recognition
[129] Data-argumentation-based Action classification, emotion recognition

### v2

@@ 3.1.2 Learning based imputation
Lin et al. proposed a novel objective by integrating representation learning and data recovery into a unified information-theoretic framework termed COMPLETER [80] for clustering task. Specifically, COMPLETER propose to learn intra-modal consistency by maximizing mutual information and learn modal imputation by minimizing conditional entropy. Dong et al. proposed a deep Gaussian processes model to impute the missing modalities and further enhance the performance of downstream tasks [81]. Motivated by contrastive learning, Yang et al. introduced a multimodal imputation paradigm termed SURE for robust multimodal imputation [82]. The method treats complete multimodal pairs as positive and randomly sampling unpaired samples as negative. A noise-robust contrastive loss is employed to alleviate the influence of false negatives. Li et al. proposed a prototype-based multimodal imputation method, which incorporates a two-stream model with dual attention layers and contrastive learning mechanism to learn modal-specific prototypes and model the cross-modal relationship [83]. Considering incomplete multimodal inputs, the model performs modality imputation using prototypes learned from the incomplete multimodal training data and the sample-prototype relationship inherited from the observed modalities. Xu et al[84] proposed a novel federated multi-view clustering method using an unsupervised technique to evaluate and refine imputation quality, efficiently handling various scenarios of incomplete multi-view data.
@@ 3.2.1 Latent representation and projection learning
Zhou et al[87] integrated the latent representation learning and regression into a unified framework for Alzheimer’s disease diagnosis based on incomplete multimodal neuroimaging and genetic data. The motivation of the latent representation learning module for incomplete multimodal data is similar to model (4). However, a widely recognized limitation of the above methods based on model (4) is that they are only applicable to the case with bi-modal data. To address this limitation, weighted matrix factorization is introduced, where doubly aligned incomplete multimodal clustering [88], one-pass incomplete multimodal clustering [89], and localized sparse incomplete multimodal clustering [90] are representative methods. These models generally impose the location information of missing instances as the weighted matrices with binary values on the matrix factorization model to eliminate the loss yet reduce the negative influence of the missing instances. The basic model of the weighted matrix factorization based methods can be formulated as follows:
@@
Compared with the basic model (4), the methods based on model (5) are more flexible and can be easily extended to the data with more than two modalities and arbitrary incomplete cases. Besides of these weighted matrix factorization models, a generalized incomplete multimodal clustering model is proposed to learn a complete common representation from the incomplete modal-specific representations with a partially aligned learning term [91]. Specifically, such a partially aligned learning term can be regarded as a variant of the weighted common representation learning. To prevent the model from overly focusing on certain modalities with larger dimensions, projective incomplete multimodal clustering transforms the matrix factorization into a variant of the weighted projection learning model to obtain the consensus representation from the incomplete multimodal data [92]. For the incomplete multimodal and incomplete multi-label classification task, a variant of the weighted projection learning model is developed to explore the information of available modalities [93]. In terms of infant neuroscience, Zhang et al. propose to explore incomplete multimodal data by constructing a latent representation to capture the complementary information from incomplete time-points [94]. In summary, the above methods focus on exploring the available information of the observed modalities by introducing some techniques like the weighting mechanism to eliminate the loss of those missing modalities.
@@ 3.2.2 Graph learning based
Graph learning based method is very popular in conventional multimodal clustering. This technique is also widely considered and integrated into other methods like matrix factorization based and deep learning based methods, to explore the structure information of data so as to enhance the performance. For incomplete multimodal learning, the main challenges to apply the graph learning based approach are two-folds: (1) How to obtain the intrinsic structure relationships among the available instances in each modality? (2) How to fully explore these incomplete relationships? For example, for the incomplete multimodal clustering tasks, one of the major challenges is to obtain the consensus complete graph or representation from these incomplete modal-specific relationships associated with all incomplete modalities. Specifically, we can formulate this issue as the following simple framework:
@@
where ${{{\bar{X}}^{\left(m\right)}}}$ denotes the set of the available instances in the $m$-th modality. ${{S^{\left(m\right)}}}$ represents the modal-specific graph learned from the $m$-th modal data ${{{\bar{X}}^{\left(m\right)}}}$. ${{W^{\left(m\right)}}}$ is the modal-available or missing indicator matrix as defined in Eq. 5. Taking the incomplete multimodal clustering task for example, the target of general graph learning based method is to learn the consensus graph or representation $Z$ from the incomplete data $\left\{{{{\bar{X}}^{\left(m\right)}}}\right\}_{m=1}^{l}$ or graph $\left\{{{S^{\left(m\right)}}}\right\}_{m=1}^{l}$, such that a consensus clustering result can be obtained by performing k-means or spectral clustering on $Z$.
As one representative work, incomplete multimodal spectral clustering [95] proposed a unified framework that integrates the adaptive collaborative representation based modal-specific graph learning and co-regularized consensus representation learning. The first term enables the method to adaptively capture the intrinsic relationships among available instances of each modality and the second term seeks to obtain the consensus complete representation by fully exploring such similarity relationship. For the data with only two modalities, Wu et al[96] proposed a method based on structured graph learning, which seeks to learn two modal-specific graphs corresponding to the two modalities with a partially aligned constraint on the samples with complete modalities. Then, the method obtains the modal-specific clustering from the two graphs and adopts the best one-to-one map to obtain the final consensus clustering result. The work in [97] pre-constructs several modal-specific graphs from the available instances of every modality and then learns a consensus graph from these graphs via a weighted consensus graph learning model with the constraint of pre-constructed confident neighbor information. In summary, except for the imputation based graph methods, most graph-based incomplete multimodal learning methods seek to exploit certain similarity information among available instances to discover the clustering structure and clustering result. In [98], for Alzheimer’s disease prediction, a unified linear regression based model is developed, in which the incomplete multimodal data are segmented into several complete subsets for model optimization. Similar to the above method, Liu et al[99] segmented the incomplete multimodal data into several complete subsets and used the sparse representation algorithm to construct the sub-graphs from these subsets. Then a modal-aligned hypergraph learning model is developed to perform the Alzheimer’s disease diagnosis.
@@ 3.2.3 Kernel learning based
Kernel learning based methods transform the learning phase from the raw-feature space into the kernel space. Similar to the graph learning based methods, each element of the kernel matrix denotes a kind of the similarity or connection values of the corresponding two samples. For incomplete multimodal data, the constructed kernels w.r.t. different modalities will also be incomplete with missing rows and columns. Thus, the target of kernel learning based methods for incomplete multimodal data is to explore the information of these incomplete kernels. In this branch, most methods consider kernel imputation as we mentioned above like model (3) while relatively few works focus on exploiting certain available kernel information. Late fusion incomplete multimodal clustering [100] is one of the representative kernel imputation-free methods for incomplete multimodal data. The method does not need to impute the missing kernel elements but tries to simultaneously impute the missing representations corresponding to the missing modalities and learn the consensus representation in a joint framework. Furthermore, to reduce the low-quality imputation of the latent representations, Liu et al[101] proposed an efficient and effective regularized incomplete multimodal clustering method (EERIMVC), which integrates the initialized consensus representation as prior knowledge. In this model, such prior representation is obtained by the conventional complete kernel algorithm with zero or average filling. Therefore, the quality of the initialized consensus representation will be an important factor to the performance. To break the gap between representation learning and label prediction, a one-stage incomplete multimodal clustering method is proposed [102]. Compared with EERIMVC, this method replaces the consensus representation with a clustering label matrix and an orthogonal matrix.
@@ 3.2.4 Deep learning based
Recent years, deep learning received lots of attention in incomplete multimodal learning. For instance, Zhao et al[103] proposed a deep partial multimodal clustering network, which first uses deep encoders to extract the latent features of multimodal data and then adopts the graph regularized partial multimodal clustering model to learn the consensus representation from the incomplete bi-modal data. For data with more than two modalities, Wen et al[104] proposed several deep incomplete multimodal clustering networks based on deep autoencoder framework, cognitive learning [105], and graph neural network [106], respectively. The objective function of these methods can be unified as the following formula:
@@
The above methods focus on the observed instances to train the network and can flexibly handle all kinds of incomplete data with more than two modalities by introducing the location matrix ${W^{\left(m\right)}}$ as prior weight in function $f$. In addition, a common technique adopted in these three methods is the weighted latent representation fusion, in which the location information of missing modalities is adopted as fusion weight to reduce the negative influence of missing modalities. Lately, Xu et al[108] proposed an imputation-free and fusion-free deep network for incomplete multimodal data clustering. The method mines the multimodal cluster complementarity in the subspace and transforms such complementary information as supervised information to guide the network training. However, this method needs a part of samples with fully observed modalities for model training.
@@ 3.3 Discussion
To sum up, the ways to handle incomplete data can be commonly divided into two categories, imputation-based incomplete multimodal learning and imputation-free incomplete multimodal learning. Methods that focus on imputing incomplete modalities can be classified into two types: one indirectly recovers the missing modalities through graph or kernel completion, and the other directly fills in the original data. For imputation-free incomplete multimodal learning methods, their design philosophy focuses more on utilizing available modalities to explore information and understand multimodal data. However, both types of methods for handling incomplete multimodal data face deeper challenges. For example, the quality assessment of imputed instances for missing modalities is commonly overlooked. Besides, using prior missing information to mask unknown modalities is itself difficult to bridge the information gap and information imbalance issues caused by modal missing.
@@ 4 Balanced Multimodal Learning
Different modalities are tightly correlated, since they depict the same concept from different views. This property inspires the flourish of multimodal learning, where multiple modalities are integrated and expected to enhance the understanding of related events or objects. However, despite the natural cross-modal correlation, each modality has distinct data sources and forms. For instance, audio data is typically represented as one-dimensional waveforms, while visual data consists of images composed of pixels. On the one hand, this discrepancy endows each modality with different properties, such as convergence speed, and then makes it hard to process and learn all modalities well simultaneously, bringing difficulty to joint multimodal learning. On the other hand, this discrepancy also reflects in the quality of unimodal data. Even though all modalities depict the same concepts, their information related to the target event or objects is in different amounts. For example, considering an audio-visual sample of label meeting, the visual data apparently show the visual content of the meeting, which is easy to recognize (see Figure.1c). While the corresponding audio data is a noisy street car sound, which is hard to build correlation with the label meeting. The information amount of visual modality is clearly more than that of audio modality. Due to the greedy nature of deep neural networks [9], multimodal model would tend to rely on just high-quality modality with sufficient target-related information, while under-fitting the other modalities. To address these challenges and improve the efficacy of multimodal models, recent studies have focused on strategies to balance the differences between modalities and enhance the overall performance of the model.
Figure 3: The learning curves (error-rate) of audio model (A), video model (V), and the naive joint audio-video (AV) model on the Kinetics dataset. Solid lines plot validation error while dashed lines show train error. Figure is from [7]
@@ 4.1 Property-discrepancy method
In most multimodal models, different modalities are commonly trained jointly with uniform learning objectives. While this joint training strategy has been generally proven to be effective in practice, there are still potential risk brought by the discrepancy in data of multiple modalities. The discrepancy of multimodal data in sources and forms result in that each modality has its own learning property. For example, as shown in Figure 3, in the action recognition dataset Kinetics [111], which contains audio and visual modalities, the audio data tends to require fewer training steps than the vision data before the model begins to be overfitting [112, 115]. The widely-used joint training multimodal framework often neglects this inherent learning property discrepancy of unimodal data, which can potentially lead to negative impacts on the model’s performance. More importantly, based on recent study [7], the jointly trained multimodal model even could be outperformed by the best unimodal counterpart. Thus, it is essential to consider the discrepancy in learning properties across each modality to improve the performance of the model. To address this issue and balance the discrepancy in learning properties, several studies have been proposed from different perspectives, including learning-objective-based method, optimization-based method and architecture-based method.
@@ 4.1.1 Learning-objective-based method
Since jointly trained multimodal frameworks often only have a uniform loss, which integrates the information of all modalities, the learning of each modality is hard to observe or control. The learning-objective-based methods add unimodal loss and make it possible to targetedly adjust unimodal training:
@@
where $\alpha_{mm}$ and $\alpha_{i}$ is the weight of loss function. $n$ is the number of modalities. Since the loss weight determines the influence of the corresponding loss function in parameter update, its selection is essential for controlling unimodal training. Wang et al. [7] suggested to determine the optimal loss weights by measuring the proposed overfitting-to-generalization ratio, by which it could comprehensively consider and balance the discrepancy in convergence speed of each modality.
@@ 4.1.3 Architecture-based method
Besides the above architecture-agnostic methods which focus on learning objective and optimization, methods considering the discrepancy in learning property problem from the perspective of model architecture are also proposed. At the fusion stage, Xiao et al. [115] proposed to randomly drop the network pathway of the modality with faster convergence speed altogether with a certain probability. In this way, the training of the corresponding modality is slowed down and its learning dynamics are ensured to be more compatible with other unimodal counterparts. Besides the late fusion architecture, studies that utilize low-level cross-modal fusion to enhance multimodal joint learning are also proposed. Zhou et al[116] utilized central connections to link low-level features in both video and audio streams, capturing spatial (for video frames) and temporal information, thereby achieving high-level semantic representations and improving joint multimodal learning.
@@
Category Application Reference
Property discrepancy Optimization-based Action classification [7]
Sentiment analysis [114]
Architecture-based Action classification [115]
Deepfake detection [116]
Quality discrepancy Learning-objective-based Violence Detection [117]
Action classification [118]
Product understanding [119]
Audio-visual fine-grained classification [120]
Action classification [121]
Optimization-based Action classification, sentiment analysis [8, 122]
Rumor detection [123]
Audio-visual video parsing [124]
Architecture-based Sentiment analysis [125]
Multimodal Conditional Image Synthesis [126]
RGB-T Semantic Segmentation [127]
Data-augmentation-based Muli-view classification, gesture classification [9]
Emotion recognition [128]
Action classification, emotion recognition [129]